At 20∘C, the hole in an aluminum ring is 2.300 cm in diameter. You need to slip this ring over a steel shaft that has a room-temperature diameter of 2.306 cm .To what common temperature should the ring and the shaft be heated so that the ring will just fit onto the shaft? Coefficients of linear thermal expansion of steel and aluminum are 12×10−6 K−1 and 23×10−6 K−1 respectively.
Answer #1

dal aal - ds as = 2.3(23 x 10^-6) - (2.306 x 12 x 10^-6) = 2.52 x 10^-5 /K

Temperature, T2 = To + dT

T2 = 20 + (2.306 - 2.3)/(2.52 x 10^-5) = 257.83 C

Comment in case any doubt please rate my answer ....